OpenStudy (anonymous):

48 What is the solution of the equation? (x-2)3/4=8 (2 points)

OpenStudy (nikato):

\[\frac{ (x-2)3 }{4 }=8 is this how \it looks?\]

OpenStudy (anonymous):

i will guess \[\large (x-2)^{\frac{3}{4}}=8\]

OpenStudy (anonymous):

neither of us answered the question, we were just trying to figure out what it said

OpenStudy (anonymous):

@LaurenEdwards"Parker" first please learn to use the equation editor to type the correct question. Or nobody can answer or help you with the correct one